A baloon is ascending at a velocity of 12 m/s A packet is dropped from it when it is at height of 65m from the ground. If packet reaches the ground after time T....then find value of T

Grade:11

2 Answers

Arun
4 years ago
initial velocity of packet = -12m/s  (taking up direction as negative)
acceleration, a = 10m/ss
height, h = 65m
thus using second equation of motion you can find out the time taken and of course the packet will reach the ground.
Saurabh Koranglekar
4 years ago
Dear student

T be t1 +t2

t1 = 12/g = 1.2 sec
Height above ground = 65 + h
h = (144/20)= 7.2 m

t2 = sqrt( 2*72.2 /g ) = 3.8

T = 5 seconds

Regards